OpenStudy (anonymous):

Find the quotient of the quantity negative 6 times x to the 2nd power times y to the 8th power plus 12 times x times y to the 3rd power minus 36 times x times y to the 2nd power all over 6 times x times y to the 2nd power. −6x2y8 + 12xy3 − 6 −xy6 + 2y − 6 xy6 + 2y − 6 −xy6 + 2xy − 6

OpenStudy (anonymous):

@johnweldon1993 help, smartiepants. :P

OpenStudy (johnweldon1993):

Sorry hun still on? Well regardless \[\large \frac{-6x^2y^8 + 12xy^3 - 36xy^2}{6xy^2}\] same thing as before...every term in the numerator is divisible by 6...and each one here has at least an 'x' and at least a 'y^2' so we can factor each of those out When we factor out a 6 we have \[\large \frac{6(-x^2y^8 + 2xy^3 - 6xy^2)}{6xy^2}\] When we then factor out an 'x' we have \[\large \frac{6x(-xy^8 + 2y^3 - 6y^2)}{6xy^2}\] and finally when we factor out a 'y^2' \[\large \frac{6xy^2(-xy^6 + 2y - 6)}{6xy^2}\] And just like last time...we now have a 6xy^2 on both the top and bottom so they cancel and we have \[\large \frac{\cancel{6xy^2}(-xy^6 + 2y - 6)}{\cancel{6xy^2}} \rightarrow -xy^6 + 2y - 6\]